OpenStudy (anonymous):

what is the equation of a circle with a center at (-2,3) and a radius of 4

OpenStudy (anonymous):

You have to plug it into the equation r^2 = (x-h)^2 + (y-K)^2 Where r is the radius and the point (h,k) being the center of the circle So it would look like: (4)^2 = [x-(-2)]^2 + [y-(3)]^2 Then simplify the equation

OpenStudy (amistre64):

explanation was good, but a little heavy on the 'doing all the work' part and not enough 'encouraging asker participation'. all in all could use some improvement.